A person can run 10 kilometers per hour in 25 minute bursts, then he must slow to 3 kilometers per hour for 5 minutes to regain energy.
Referring to the above scenario, how many kilometers (round to the nearest integer) can this person cover in five hours?

Answer by Boreal(15235)

Every half hour, the person runs 10km/h*(25/60)h=250/60=4 1/6 km PLUS 3km/h(1/12)h=0.25 km. That sum is 4 1/6+1/4=4 5/12=53/12 km
That is done 10 times in 5 hours or 530/12 km=44 km to the nearest integer.